OpenStudy (anonymous):

What is the simplified form of (9x^2-25)/(3x+5)? choices: 3x + 5, with the restriction x ≠ 5/3 3x – 5, with the restriction x ≠ 5/3 3x – 5, with the restriction x ≠ – 5/3 3x + 5, with the restriction x ≠ –5/3

OpenStudy (anonymous):

I'm stuck between these two choices: 3x – 5, with the restriction x ≠ 5/3 3x – 5, with the restriction x ≠ – 5/3

OpenStudy (anonymous):

Medal will be given.

OpenStudy (whpalmer4):

Which of those causes 3x-5 to equal 0?

OpenStudy (anonymous):

3x – 5, with the restriction x ≠ – 5/3?

OpenStudy (whpalmer4):

let's see: 3(-5/3) - 5 = -15/3 - 5 = -5 -5 =

OpenStudy (whpalmer4):

Just set the denominator equal to 0 and solve for \(x\) \[3x-5=0\]\[3x=5\]\[x=\]

OpenStudy (anonymous):

Oh I see! Thank you so much!!